# Complexity of Finite Borel Asymptotic Dimension

### v1

@@ 1. Introduction
The study of Borel graphs has figured prominently in descriptive set theory over the past several decades (see, e.g., [KST99], [KM04], [Jac19], [Má19], [KM20], [Pik21], [Ber22], [GV24], [Kec24]). One of the central themes in this area is the notion of Borel hyperfiniteness. Recall that a Borel graph $\mathcal{G}$ is Borel hyperfinite if it can be written as a countable increasing union of Borel graphs with finite connected components, i.e., $\mathcal{G}=\bigcup_{n\in\mathbb{N}}\mathcal{G}_{n}$, where for each $n\in\mathbb{N}$, $\mathcal{G}_{n}$ is a Borel graph with finite connected components.
Hyperfinite structures arise in various areas of mathematics and theoretical computer science as non-trivial structures of minimal complexity (see, e.g., [Con76, CFW81, HKL90, KM04, Mat12, Lov12, Ele18]). The combinatorial properties of these structures often reflect the combinatorial properties of their finite substructures, for instance in the theory of graph limits [Lov12]. However, Borel hyperfiniteness is in general not strong enough to guarantee that the Borel combinatorics of a Borel graph reflects its classical combinatorics. For instance, the classical theorem of Brooks, which states that graphs of maximum degree $d\in\mathbb{N}$ have chromatic number bounded by $d$, fails in the Borel context [Mar16, Theorem 1.3], even for hyperfinite Borel graphs [CJM+20, Theorem 1.4].
In [CJM+23], Conley, Jackson, Marks, Seward, and Tucker-Drob introduce the notion of Borel asymptotic dimension as an extension of Gromov’s classical notion of asymptotic dimension [Gro93] to the Borel context. When a locally finite Borel graph $\mathcal{G}$ has finite Borel asymptotic dimension, then not only is it Borel hyperfinite [CJM+23, Theorem 7.1], but its Borel combinatorics also aligns more closely with its classical combinatorics; for instance, the Borel chromatic number of $\mathcal{G}$ is bounded above by $12\chi(\mathcal{G})-$ 1, where $\chi(\mathcal{G})$ is the classical chromatic number of $\mathcal{G}$ [CJM+23, Theorem 8.1].
It is immediate from the definitions that both finite Borel asymptotic dimension and Borel hyperfiniteness are $\mathbf{\Sigma}^{1}_{2}$ notions. While determining the exact complexity of Borel hyperfiniteness remains a major open problem in descriptive set theory [DJK94], our main result shows that finite Borel asymptotic dimension is $\mathbf{\Sigma}^{1}_{2}$-complete.
@@
Theorem 1.1 builds on the celebrated paper of Todorčević and Vidnyánszky [TV21], who showed that locally finite Borel graphs having finite Borel chromatic number form a $\mathbf{\Sigma}^{1}_{2}$-complete set, and adds to the growing body of complexity results in Borel combinatorics (see, e.g., [BCG+21], [Tho22], [FSV24]).
The main tool we use to prove Theorem 1.1 involves transforming the geometric question of whether a graph generated by a single Borel function has finite Borel asymptotic dimension (see Section 5 for the definition) into a purely combinatorial question of the existence of certain forward-recurrent sets. A Borel function $f$ on a standard Borel space $X$ is acyclic if $f^{k}(x)\neq x$ for every $x\in X$ and $k\in\mathbb{N}^{+}=\mathbb{N}\setminus\{0\}$. We define $\mathcal{G}_{f}$ to be the Borel graph on $X$ generated by $f$, that is, the graph where vertices $x,y\in X$ are adjacent if and only if $x\neq y$ and either $f(x)=y$ or $f(y)=x$. A set $H\subseteq X$ is hitting for $f$ if, for each $x\in X$, there is $k\in\mathbb{N}^{+}$ such that $f^{k}(x)\in H$. Given $r\in\mathbb{N}^{+}$, the set $H$ is $r$-forward-independent for $f$ if $f^{k}(x)\in H$ implies that $k>r$ for any $x\in H$ and $k\in\mathbb{N}^{+}$.
@@
If $f$ is bounded-to-one, then (iii) in Theorem 1.2 is always satisfied [KST99, Corollary 5.3]. We leave as an open problem whether Theorem 1.1 holds for bounded-degree Borel graphs; see Section 6.
As an application of the techniques used to prove Theorems 1.1 and 1.2, we derive a new complexity result for a homomorphism problem for the class of Borel directed graphs (digraphs for short) generated by a single Borel function. Given a finite digraph $H$, we follow Thornton [^1] We refer the reader to [Tho22] for a thorough discussion of the motivation for and connection between classical and Borel constraint satisfaction problems (CSPs). We remark that Feder and Vardi [FV99] proved that every fixed template CSP is polynomial-time equivalent to a digraph CSP. [Tho22] and write $\operatorname{CSP}_{B}(H)$ for the set (of codes) of Borel digraphs that admit a Borel homomorphism to $H$, and define $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$ to be the set (of codes) of Borel digraphs generated by a single Borel function that admit a Borel homomorphism to $H$. Note that when investigating $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$, we can without loss of generality assume that $H$ is sinkless, i.e., that every vertex of $H$ has at least one outgoing edge.
Recall that a finite digraph $H$ is ergodic if there are $v\in V(H)$ and $k_{0}\in\mathbb{N}$ such that, for every $k\geq k_{0}$, there is a directed path of length $k$ that starts and ends at $v$. A loop is a directed edge of the form $(v,v)$ for some $v\in V(H)$. Note that any digraph containing a loop is ergodic.
@@
To be more specific, recall first that any locally checkable labeling (LCL) problem on oriented paths can be thought of as a digraph homomorphism problem $\Pi_{H}$ to a finite sinkless digraph $H$ [BHK+17]. The classification of LCL problems on oriented paths in the LOCAL model [BHK+17] states that:
@@
    If $H$ contains a loop, then $\Pi_{H}$ can be solved in 00 rounds of communication in the LOCAL model.
@@
Figure 1. (a) A directed graph $H$ such that $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$ is $\mathbf{\Pi}^{1}_{1}$. (b) An abstract walk $p$. (c) The directed graph $H^{p}$, which has the property that $\operatorname{CSP}^{\operatorname{function}}_{B}(H^{p})$ is $\mathbf{\Sigma}^{1}_{2}$-complete, showing that $\operatorname{CSP}_{B}(H)$ is $\mathbf{\Sigma}^{1}_{2}$-complete as well.
Remark 1.5.
As $\mathbf{\Sigma}^{1}_{2}$ is an upper bound on the complexity of $\operatorname{CSP}_{B}(H)$, it follows that if $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$ is $\mathbf{\Sigma}^{1}_{2}$-complete, then so is $\operatorname{CSP}_{B}(H)$. The reverse implication is easily seen to fail.
Suppose now that $H$ is both sinkless and sourceless [^2] The finite and Borel CSP dichotomy coincide in this case [BKN09, Tho22]. Let $p$ be an abstract walk, i.e., a finite sequence of arrows as in Figure 1(b). Define $H^{p}$ to be the $p$-power of $H$, i.e., the directed graph on $V(H)$ where there is a directed edge from $x$ to $y$ if there is a realization in $H$ of the walk $p$ from $x$ to $y$, see Figure 1(c) for an example. Since $H$ is sinkless and sourceless, it can be easily seen that $H^{p}$ is also sinkless and sourceless. If $\operatorname{CSP}^{\operatorname{function}}_{B}(H^{p})$ is $\mathbf{\Sigma}^{1}_{2}$-complete for some walk $p$, then so is $\operatorname{CSP}_{B}(H)$. We leave as an open problem whether the converse holds; see Section 6.
The paper is organized as follows. In Section 2, we provide the requisite preliminaries. In Section 3, we prove the main complexity result that we use throughout the rest of the paper. The proof of Theorem 1.4 is given in Section 4. In Section 5, we prove Theorems 1.1 and 1.2. We conclude with some open questions in Section 6.
@@ 2. Preliminaries
Recall that $[\mathbb{N}]^{\mathbb{N}}$ denotes the space of infinite subsets of $\mathbb{N}$. Each point $x\in[\mathbb{N}]^{\mathbb{N}}$ may be identified with its increasing enumeration $x=(x(n))_{n\in\mathbb{N}}$. The shift function $S:[\mathbb{N}]^{\mathbb{N}}\to[\mathbb{N}]^{\mathbb{N}}$ is defined by
@@
and the shift graph $\mathcal{G}_{S}$ is the (undirected) graph on $[\mathbb{N}]^{\mathbb{N}}$ generated by $S$.
@@
Let $R$ be a relation on $X$, let $S$ be a relation on a standard Borel space $Y$, and assume that $R$ and $S$ have the same arity $d\in\mathbb{N}$. A Borel map $\varphi:X\to Y$ is a homomorphism from $(X,R)$ to $(Y,S)$, or from $R$ to $S$ for short, if
@@
A finite digraph $H=(V(H),E(H))$ is sinkless if every vertex has an outgoing edge and sourceless if every vertex has an ingoing edge. A directed path in $H$ is a sequence of vertices $(v_{0},\dots v_{k})\subseteq V(H)$ such that $(v_{i},v_{i+1})\in E(H)$ for every $10\leq i\leq k-$ 1. A strong connectivity component of $H$ is an inclusion-maximal set $A\subseteq V(H)$ such that, for any vertices $v,w\in A$, there is a directed path from $v$ to $w$ in $H\restriction A$. We say that $H$ is strongly connected if $V(H)$ is a strong connectivity component. Note that every finite sinkless digraph contains at least one strong connectivity component since it contains some vertex $v$ for which there is a directed path from $v$ to itself.

### v2

@@ 1. Introduction
The study of Borel graphs and Borel equivalence relations has figured prominently in descriptive set theory over the past several decades (see, e.g., [28, 27, 25, 33, 26, 38, 4, 20, 30]). One of the central themes in this area is the notion of Borel hyperfiniteness. A Borel graph $\mathcal{G}$ is Borel hyperfinite if it can be written as a countable increasing union of Borel graphs with finite connected components, i.e., $\mathcal{G}=\bigcup_{n\in\mathbb{N}}\mathcal{G}_{n}$, where for each $n\in\mathbb{N}$, $\mathcal{G}_{n}$ is a Borel graph with finite connected components and $\mathcal{G}_{n}\subseteq\mathcal{G}_{n+1}$.
Hyperfinite structures arise in various areas of mathematics and theoretical computer science as non-trivial structures of minimal complexity (see, e.g., [12, 11, 22, 27, 34, 31, 16]). In particular, in the theory of countable Borel equivalence relations, the class of Borel hyperfinite equivalence relations is the minimum non-trivial element in the Borel reducibility hierarchy by the Glimm–Effros dichotomy [19, 14, 15, 22]. However, many basic problems about Borel hyperfiniteness remain open (see [30] for a recent survey). Among these problems, one of the deepest centers around the complexity of hyperfiniteness: While it is immediate from the definition that hyperfiniteness is a $\mathbf{\Sigma}^{1}_{2}$ notion, it is unknown whether the set of codes of Borel hyperfinite graphs is $\mathbf{\Sigma}^{1}_{2}$-complete, that is, maximally complex [13] (see Section 2 for the formal definitions of the complexity notions). The $\mathbf{\Sigma}^{1}_{2}$-completeness of Borel hyperfiniteness would rule out a number of conjectured dichotomy theorems [30].
In [10], the authors introduce the notion of Borel asymptotic dimension as an extension of Gromov’s classical notion of asymptotic dimension for metric spaces [21] to the definable context; they also show that if a locally finite Borel graph has finite Borel asymptotic dimension, then it is Borel hyperfinite [10, Theorem 7.1]. Despite its recency, Borel asymptotic dimension has become an invaluable tool in the study of Borel hyperfiniteness (see, e.g., [3, 36, 24]).
While determining the exact complexity of Borel hyperfiniteness remains a major open problem, our main result shows that the notion of finite Borel asymptotic dimension is $\mathbf{\Sigma}^{1}_{2}$-complete.
@@
Theorem 1.1 builds on the celebrated paper of Todorčević and Vidnyánszky [41], who showed that locally finite Borel graphs having finite Borel chromatic number form a $\mathbf{\Sigma}^{1}_{2}$-complete set, and adds to the growing body of complexity results in Borel combinatorics (see, e.g., [6, 40, 18]).
The main tool we use to prove Theorem 1.1 involves transforming the geometric question of whether a graph generated by a single Borel function has finite Borel asymptotic dimension (see Section 5 for the definition) into a purely combinatorial question concerning the existence of certain forward-recurrent sets. A Borel function $f$ on a standard Borel space $X$ is acyclic if $f^{k}(x)\neq x$ for every $x\in X$ and $k\in\mathbb{N}^{+}=\mathbb{N}\setminus\{0\}$. We write $\mathcal{G}_{f}$ for the Borel graph generated by $f$, that is, the graph on $X$ where vertices $x,y\in X$ are adjacent if and only if $x\neq y$ and either $f(x)=y$ or $f(y)=x$. A set $H\subseteq X$ is hitting for $f$ if, for each $x\in X$, there is $k\in\mathbb{N}^{+}$ such that $f^{k}(x)\in H$. Given $r\in\mathbb{N}^{+}$, the set $H$ is $r$-forward-independent for $f$ if $f^{k}(x)\in H$ implies that $k>r$ for any $x\in H$ and $k\in\mathbb{N}^{+}$.
@@
    For each $r\in\mathbb{N}^{+}$, there is a Borel $r$-forward-independent hitting set for $f$.
@@
Condition (iii) in Theorem 1.2 is satisfied if, for instance, $f$ is bounded-to-one. This follows from [28, Corollary 5.3]. Furthermore, (iii) is also satisfied when the connectedness relation of $\mathcal{G}_{f}$ admits a Borel transversal.
Borel graphs with finite Borel asymptotic dimension enjoy certain definable combinatorial properties that are not shared by Borel hyperfinite graphs in general. For instance, Brooks’s theorem, a fundamental result in the theory of graph colorings, fails for Borel colorings [32], even when the graph under consideration is Borel hyperfinite [9]. However, a version of Brooks’s theorem for Borel colorings, and in fact Borel versions of many local coloring problems that in general only admit solutions up to null sets, can be recovered for Borel graphs with finite Borel asymptotic dimension [2].
As an application of the techniques used to prove Theorems 1.1 and 1.2, we derive a new complexity result for a homomorphism problem for the class of Borel directed graphs (digraphs for short) generated by a single Borel function. Given a finite digraph $H$, we follow Thornton [40] and write $\operatorname{CSP}_{B}(H)$ for the set (of codes) of Borel digraphs that admit a Borel homomorphism to $H$, and we also define $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$ to be the set (of codes) of Borel digraphs generated by a single Borel function that admit a Borel homomorphism to $H$. Note that when investigating $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$, we can without loss of generality assume that $H$ is sinkless, i.e., that every vertex of $H$ has at least one outgoing edge.
Recall that a finite digraph $H$ is ergodic if there are $v\in V(H)$ and $k_{0}\in\mathbb{N}$ such that, for every $k\geq k_{0}$, there is a directed path of length $k$ that starts and ends at vv. [^1] Here, the motivation for the term “ergodic” originates in the theory of Markov chains, where a transition digraph is said to be ergodic if it is irreducible and aperiodic, see, for example, [37]. In our context, the term is used for digraphs that contain a strongly connected component that is ergodic in the stochastic sense. A loop is a directed edge of the form $(v,v)$ for some $v\in V(H)$. Note that any digraph containing a loop is ergodic.
@@
Figure 1. (a) A directed graph $H$ such that $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$ is $\mathbf{\Pi}^{1}_{1}$. (b) An abstract walk pp. (c) The directed graph $H^{p}$, which has the property that $\operatorname{CSP}^{\operatorname{function}}_{B}(H^{p})$ is $\mathbf{\Sigma}^{1}_{2}$-complete, showing that $\operatorname{CSP}_{B}(H)$ is $\mathbf{\Sigma}^{1}_{2}$-complete as well
Remark 1.5 (Connections to CSPs).
Theorem 1.4 is motivated by the classical theory of constraint satisfaction problems (CSPs); see, for example, [5] for a comprehensive overview of the subject. A (fixed-template) constraint satisfaction problem takes the following form: Given a finite structure $H$, how complicated is it to decide whether a finite structure in the same signature admits a homomorphism to $H$? The target structure $H$ is called the template, and the set of finite structures admitting a homomorphism to $H$ is denoted $\operatorname{CSP}(H)$. The well-known CSP dichotomy conjecture, now a theorem, states that every CSP is either solvable in polynomial time or NP-complete.
In their foundational paper, Feder and Vardi [17] showed that every CSP is polynomial-time equivalent to a digraph CSP, i.e., a CSP whose template is a digraph; therefore, resolving the full conjecture amounts to resolving it just for digraph CSPs. A special case of the CSP dichotomy in the case when the template is a sinkless and sourceless digraph was established by Barto, Kozik, and Niven [1]. Later, the full CSP dichotomy conjecture was confirmed independently by Bulatov [8] and by Zhuk [42, 43].
Recently, Thornton [40] initiated the study of the complexity of CSPs in measurable combinatorics. More precisely, given a finite digraph $H$, we may consider, as above, the set $\operatorname{CSP}_{B}(H)$ (of codes) of Borel digraphs that admit a Borel homomorphism to $H$. Among other results, Thornton [40] proved a Borel analogue of the CSP dichotomy for sinkless and sourceless digraphs; in particular, when $H$ is sinkless and sourceless, $\operatorname{CSP}_{B}(H)$ is either $\mathbf{\Pi}^{1}_{1}$ or $\mathbf{\Sigma}^{1}_{2}$-complete. Moreover, these two cases correspond exactly to the classical ones: $\operatorname{CSP}_{B}(H)$ is $\mathbf{\Pi}^{1}_{1}$ when $\operatorname{CSP}(H)$ is solvable in polynomial time, and $\operatorname{CSP}_{B}(H)$ is $\mathbf{\Sigma}^{1}_{2}$-complete when $\operatorname{CSP}(H)$ is NP-complete.
Remark 1.6 ($\operatorname{CSP}_{B}(H)$ vs $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$).
Note that, when working with $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$, we do not impose restrictions on the templates themselves; instead, we consider a subclass of digraphs on which the existence of Borel homomorphisms into $H$ is queried. For any digraph $H$, since $\mathbf{\Sigma}^{1}_{2}$ is an upper bound for the complexity of $\operatorname{CSP}_{B}(H)$, if $\operatorname{CSP}^{\operatorname{function}}_{B}(H)$ is $\mathbf{\Sigma}^{1}_{2}$-complete, then $\operatorname{CSP}_{B}(H)$ is $\mathbf{\Sigma}^{1}_{2}$-complete as well. The converse does not hold in general (see Figure 1 and its description for a counterexample). However, it is possible that a weaker form of the converse holds; see the next paragraph and Section 6.
Suppose now that $H$ is both sinkless and sourceless. Let $p$ be an abstract walk, i.e., a finite sequence of arrows as in Figure 1(b). Define $H^{p}$ to be the $p$-power of $H$, i.e., the directed graph on $V(H)$ where there is a directed edge from $x$ to $y$ if there is a realization in $H$ of the walk $p$ from $x$ to $y$ (see Figure 1(c) for an example). Since $H$ is sinkless and sourceless, it can be easily seen that $H^{p}$ is also sinkless and sourceless. If $\operatorname{CSP}^{\operatorname{function}}_{B}(H^{p})$ is $\mathbf{\Sigma}^{1}_{2}$-complete for some walk $p$, then so is $\operatorname{CSP}_{B}(H)$; this follows from the fact that, for any function $f$, the directed graph $\overrightarrow{\mathcal{G}}_{f}$ generated by $f$ admits a (Borel) homomorphism to $H^{p}$ if and only if the directed graph formed from $\overrightarrow{\mathcal{G}}_{f}$ by replacing each directed edge with the path $p$ admits a (Borel) homomorphism to $H$ (we note that this construction is definable in the codes). We leave as an open problem whether the converse holds; see Section 6.
Remark 1.7 (Connections to the LOCAL model).
@@
To be precise, recall first that any locally checkable labeling (LCL) problem on oriented paths can be thought of as a digraph homomorphism problem $\Pi_{H}$ to a finite sinkless digraph $H$ [7]. The classification of LCL problems on oriented paths in the LOCAL model [7] states that:
@@
    If $H$ contains a loop, then $\Pi_{H}$ can be solved in 0 rounds of communication in the LOCAL model.
@@
The paper is organized as follows. In Section 2, we provide some preliminaries. In Section 3, we prove the main complexity result that we use throughout the rest of the paper. The proof of Theorem 1.4 is given in Section 4. In Section 5, we prove Theorems 1.1 and 1.2. We conclude with some open questions in Section 6.
@@ 2. Preliminaries
Recall that $\mathbb{N}$$^{\mathbb{N}}$ denotes the space of infinite subsets of $\mathbb{N}$. Each point $x\in[\mathbb{N}]^{\mathbb{N}}$ may be identified with its increasing enumeration $x=(x(n))_{n\in\mathbb{N}}$. The shift function $S:[\mathbb{N}]^{\mathbb{N}}\to[\mathbb{N}]^{\mathbb{N}}$ is defined by
@@
and the shift graph $\mathcal{G}_{S}$ is the (undirected) graph on $\mathbb{N}$$^{\mathbb{N}}$ generated by $S$.
@@
Let $R$ be a Borel relation on $X$, let $S$ be a Borel relation on a standard Borel space $Y$, and assume that $R$ and $S$ have the same arity $d\in\mathbb{N}$. A Borel map $\varphi:X\to Y$ is a homomorphism from $(X,R)$ to $(Y,S)$, or from $R$ to $S$ for short, if
@@
A finite digraph $H=(V(H),E(H))$ is sinkless if every vertex has an outgoing edge and sourceless if every vertex has an ingoing edge. A directed path in $H$ is a sequence of vertices $(v_{0},\dots v_{k})\subseteq V(H)$ such that $(v_{i},v_{i+1})\in E(H)$ for every $0\leq i\leq k-1$. A strong connectivity component of $H$ is an inclusion-maximal set $A\subseteq V(H)$ such that, for any vertices $v,w\in A$, there is a directed path from $v$ to $w$ in $H\restriction A$. We say that $H$ is strongly connected if $V(H)$ is a strong connectivity component. Note that every finite sinkless digraph contains at least one strong connectivity component since it contains some vertex $v$ for which there is a directed path from $v$ to itself.
${\bf\Sigma}^{1}_{2}$-completeness. The projective hierarchy provides a general framework for the study of complexity problems in measurable combinatorics. Recall that, given a Polish space $X$, a set $A\subseteq X$ is $\mathbf{\Sigma}^{1}_{2}$ if it is the projection of a ${\bf\Pi}^{1}_{1}$ (co-analytic) set (see [29, Chapter V]). For example, it is easy to see that the set of (codes of) locally finite Borel graphs having finite Borel chromatic number is $\mathbf{\Sigma}^{1}_{2}$. More generally, given any local graph coloring problem, the set of (codes of) locally finite Borel graphs admitting a solution to the problem is $\mathbf{\Sigma}^{1}_{2}$. We refer the reader to [35] and [18, Section 1] for an overview of the coding method.
Definition 2.1 (${\bf\Sigma}^{1}_{2}$-completeness).
Let $A$ be a subset of a Polish space $X$. We say that $A$ is ${\bf\Sigma}^{1}_{2}$-hard if for every Polish space $Y$ and every ${\bf\Sigma}^{1}_{2}$ set $B\subseteq Y$, there is a Borel map $f:Y\to X$ such that $f^{-1}(A)=B$, i.e., $f(y)\in A$ if and only if $y\in B$. A set $A\subseteq X$ is called ${\bf\Sigma}^{1}_{2}$-complete if it is simultaneously ${\bf\Sigma}^{1}_{2}$-hard and ${\bf\Sigma}^{1}_{2}$.
As mentioned in the introduction, Todorčević and Vidnyánszky [41] showed that the set of codes of locally finite Borel graphs that have finite Borel chromatic number is a ${\bf\Sigma}^{1}_{2}$-complete set. In our arguments, we use a general result for proving ${\bf\Sigma}^{1}_{2}$-completeness that builds on [41] and was derived recently by Frisch, Shinko and Vidnyánszky [18, Theorem 3.1]. We refer the reader to [18, Section 1] for the definition of Borel $L$-structure.
Theorem 2.2 (Theorem 3.1, [18]).
Let $\mathcal{H}$ be a Borel $L$-structure on some Polish space $Z$, and assume that there exist a Borel $L$-structure $\mathcal{G}$ on $\mathbb{N}$$^{\mathbb{N}}$ that does not admit a Borel homomorphism to $\mathcal{H}$ and a Borel map $\Phi:\text{Dom}\to Z$ so that, for each $x$, we have that $\Phi_{x}$ is a homomorphism from $\mathcal{G}\upharpoonright\text{Dom}_{x}$ to $\mathcal{H}$. Then the Borel $L$-structures that admit a Borel homomorphism to $\mathcal{H}$ form a ${\bf\Sigma}^{1}_{2}$-complete set.
In fact, we use a slight strengthening of Theorem 2.2 that follows from the proof of [18, Theorem 3.1]. Given an $L$-structure $\mathcal{G}$ on $\mathbb{N}$$^{\mathbb{N}}$, write $\mathcal{G}^{\prime}$ for the $L$-structure on $\mathbb{N}^{\mathbb{N}}\times[\mathbb{N}]^{\mathbb{N}}$ where each vertical section is a copy of $\mathcal{G}$ and no additional elements are related.
Theorem 2.3.
Let $\mathcal{H}$ be as in Theorem $2.$ Then there is a Borel set $B\subseteq\mathbb{N}^{\mathbb{N}}\times\mathbb{N}^{\mathbb{N}}\times[\mathbb{N}]^{\mathbb{N}}$ such that the set
[…]
is ${\bf\Sigma}^{1}_{2}$-complete.